OpenStudy (anonymous):

An airplane is flying in a horizontal circle at a speed of 104 m/s. The 86.0 kg pilot does not want the centripetal acceleration to exceed 6.49 times free-fall acceleration. Find the minimum radius of the plane’s path. The acceleration due to gravity is 9.81 m/s2 at this radius, what is the net force that maintain circular motion exerted on the pilot by the seat belts, the friction against the seat and so forth?

OpenStudy (loser66):

we have formula for centripetal acceleration \(\alpha = \dfrac{v^2}{r}\) where \(v= 104m/s\) and \(\alpha = 6.49 g= 6.49*9.8 = 63.602\) now replace to \(r = \dfrac{v^2}{\alpha}\) we have r = \(\dfrac{104^2}{63.602}= 170.1m\)

OpenStudy (loser66):

at this radius, the net force that maintain circular motion is centripetal force which is calculated by F = \(\dfrac{mv^2}{r} = \dfrac{86.0*104^2}{170.1}=5.4*10^3 N\)